# Accelerated Parallel Tempering via Neural Transports

### v2

@@ A.4 Parallelism versus Acceleration Time
Given two distributions $\pi^{n-1}$ and $\pi^{n}$, should we apply $K$-step forward and backward kernels; or insert $K-1$ intermediate distributions $(\pi^{n-1,k})_{k=1}^{K-1}$ and use only one-step forward and backward kernels instead?
By Proposition 2, the inverse of the local round trip rate between $\pi^{n-1}$ and $\pi^{n}$ for the first method (time-accelerated) is
@@
The following proposition analyses these local rates for both methods. As in Section A.3, the main idea is to find a vanilla PT equivalent for both algorithms. We define
@@
This proposition shows that it is preferable to use the parallel-accelerated method, as the quantity in Equation 17 is generally greater than that of Equation 16 thanks to the effect of the bridge between ${\mathbb{S}}_{0}$ and ${\mathbb{S}}_{K}$. However, in practice the time-accelerated method consumes less memory and so might be more suitable in certain circumstances.
@@ Appendix B Further Details on Analysis of PT
B.1 Proof of Proposition 2
Let $S^{n}_{t}$ be the indicator random variable with $S^{n}_{t}=1$ if the $n$-th swap is proposed and accepted at iteration $t$, and $S^{n}_{t}=0$ otherwise. By 1, the random variables $S^{n}_{t}$ are i.i.d. in $t$ with $S^{n}_{1},S^{n}_{2},\ldots\stackrel{{\scriptstyle d}}{{=}}\mathrm{Bernoulli}(s_{n})$, where $s_{n}:=1-r({\mathbb{P}}^{n-1},{\mathbb{Q}}^{n})$. This implies that the index processes $(I^{m}_{t},\epsilon^{m}_{t})$ form Markov chains on $\{0,\ldots,N\}\times\{-1,1\}$ with initial conditions $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ when $m$ is even, $\epsilon^{m}_{0}=1$ when $m$ is odd. For $t\geq 1$, each process satisfies the following recursion:
@@ B.1 Proof of Proposition 2
$\epsilon^{m}_{t}$ =$=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$ (19)
The remainder of the proof follows identically to [Syed et al., 2022, Corollary 2].
B.2 Proof of Proposition 3
We first spell out the full statement of Proposition 3.
@@ B.2 Proof of Proposition 3
where the second integral is integrated backwards in time. Assume that $f(t,\cdot)$ and $f(\cdot,x)$ are Lipschitz for all $t$ and $x$ with a global constant; and that the same holds for $b$. Moreover, suppose that there exists a $G\geq 0$ such that $||f(t,x)||+||b(t,x)||\leq G(1+||x||)$. Then
@@
where $r(p,q)=\|p\times q-q\times p\|_{\mathrm{TV}}$ and ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ are respectively the full path measures of Equation 20 and Equation 21.
@@
where the first inequality is elementary and the second follows from Proposition 6. By data processing inequality
@@
Equation 24 and Equation 25 establish Equation 23. On the other hand, Fatou’s lemma gives
$\begin{split}\liminf_{K\to\infty}r({\mathbb{P}}_{\infty|K}^{n-1},{\mathbb{Q}}_{\infty|K}^{n})&=\liminf_{K\to\infty}\frac{1}{2}\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}\left[\left|\frac{\mathrm{d}({\mathbb{Q}}_{\infty|K}^{n}\times{\mathbb{P}}_{\infty|K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n})}-1\right|\right]\\ &=\liminf_{K\to\infty}\frac{1}{2}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}\left[\left|\frac{\mathrm{d}({\mathbb{Q}}_{\infty|K}^{n}\times{\mathbb{P}}_{\infty|K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n})}-1\right|\right]\\ &\geq\frac{1}{2}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}\left[\liminf_{K\to\infty}\left|\frac{\mathrm{d}({\mathbb{Q}}_{\infty|K}^{n}\times{\mathbb{P}}_{\infty|K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n})}-1\right|\right]\\ &=\frac{1}{2}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}\left[\left|\frac{\mathrm{d}({\mathbb{Q}}_{\infty}^{n}\times{\mathbb{P}}_{\infty}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n})}-1\right|\right]=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})\end{split}$ (26)
where the penultimate equality follows from the pathwise convergence of the Euler-Maruyama schemes (Kloeden and Neuenkirch [2007], Gyöngy [1998], see also Berner et al. [2025, Lemma B.7.a]). Equation 25 and Equation 26 entail Equation 22, finishing the proof. ∎
@@
Consider the continuous-time extension $(\hat{X}_{t})_{t\in[0,1]}$ of the $K$-step Euler-Maruyama discretization of the SDE in Equation 27 given by
@@
We use ^=^$\hat{{\mathbb{P}}}_{\infty}=\hat{{\mathbb{P}}}^{(K)}_{\infty}$ to denote its path-measure. By Pinsker’s inequality
@@
[…]
@@
[…]
@@ B.3.1 Regularity Assumptions
[…]
@@
We will make the following regularity assumptions on ${\mathbb{P}}^{\bm{\beta}}$ and ${\mathbb{Q}}^{\bm{\beta}}$ analogous to Assumptions 5–7 in Syed et al. [2024]. Suppose there exists $\mathcal{F}=(\mathcal{F}_{i})_{i\in\mathbb{N}}$ of a nested sequence $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots$ of vector spaces of measurable function $f:\mathcal{X}^{K+1}\to\mathbb{R}$, such that (1) $\mathcal{F}_{i}$ contains the constant functions, and (2) $\mathcal{F}_{i}$ is closed under domination, i.e., if $g$ is a measurable function such that $|g|\leq|f|$ for some $f\in\mathcal{F}_{i}$, then $g\in\mathcal{F}_{i}$.
@@
B.3.2 Proof of Theorem 2
[…]
@@ B.3.2 Proof of Theorem 2
[…]
@@
Given an annealing schedule $10=\beta_{0}<\cdots<\beta_{N}=$ 1, let $\bm{\beta}_{n}=(\beta_{n-1},\beta_{n})$ and $\Delta\beta_{n}=\beta_{n}-\beta_{n-1}$. It follows from Lemma 3 that the sum of the rejection satisfies,
@@
We will first approximate the first term in Equation 28. Note that since $\Delta W^{\beta}=0$ and $\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}=\lim_{\beta^{\prime}\to\beta}\Delta W^{\bm{\beta}}/\Delta\beta=:\Delta\dot{W}^{\bm{\beta}}$, we have that $\dot{\alpha}^{\beta}$ satisfies:
@@
For the second term in Equation 28, we have
@@
where $\bar{W}_{2}\oplus\bar{W}_{2}(x_{0:K},x_{0:K}^{\prime}):=\bar{W}_{2}(x_{0:K})+\bar{W}_{2}(x_{0:K}^{\prime})$. 2 guarantees that the expectation in the second-to-last line is continuous in $\bm{\beta}$, and hence $C_{2}$ is finite. Next, we note that since $|\Delta\dot{W}^{\beta}(x_{0:K},x_{0:K}^{\prime})|=|\Delta\dot{W}^{\beta}(x_{0:K}^{\prime},x_{0:K})|$ is symmetric and ${\mathbb{P}}^{\beta}={\mathbb{Q}}^{\beta}$, we have:
@@
By Lemma 2, $\beta\mapsto\lambda_{\beta}$ is continuously differentiable. Since $\lambda_{\beta}\Delta\beta$ is a right Riemann sum for the integral of $\lambda_{\beta}$ with error:
@@ C.1.1 Work Formula for Deterministic flows
then, for all $(z^{n-1},z^{n})$ such that $T^{n}(z^{n-1})=z^{n}$, we have the following expression for the work defined in Equation 3:
@@
and the acceptance rate $\alpha^{n}$ defined in Section 3.2 becomes
@@
Similarly, applying Equation 30 at $(z^{n-1},z^{n})=(x^{n-1}_{*},x^{n})\in S$ gives
@@
Together Equation 31 and Equation 32 establish the proposition. ∎
@@ C.1.2 Training
Since APT provides approximate samples from both the target and reference densities for each flow, it enables a range of training objectives. Some choices include maximum likelihood estimation (MLE) (equivalent to forward KL), reverse KL, and symmetric KL (SKL), which averages the two. Each has trade-offs: forward KL promotes mode-covering, reverse KL is more mode-seeking, and SKL balances both. APT’s parallel structure makes SKL particularly effective by providing access to samples at each intermediate annealing distribution, a feature many other methods lack. For example, sequential Monte Carlo (SMC)-based approaches such as FAB [Midgley et al., 2023] and CRAFT [Matthews et al., 2022] rely on samples from only one side, limiting their choice of loss functions.
Additionally, one can explore loss functions based on APT’s rejection rates, such as the analytic round trip rate, see Proposition 2. Since higher round trip rates indicate more efficient mixing, optimizing for this metric improves sampling performance. Empirically, we found that using SKL yielded the most stable and robust results across different settings. This loss was therefore used in our final experiments. However, we leave it to future work to study other possible losses in more detail.
There are also possible variants to the training pipeline. In each case, we initialize normalizing flows to the identity transformation. One option is to run the APT algorithm with the current flows. After every (or several) steps of the APT algorithm, the current samples can be used to update the flow parameters. Since the flows are initialized at the identity transformation, the initial sampling of APT behaves similarly to PT. A second possible training pipeline is to instead use PT directly to generate a large batch of samples, and then use this batch of samples to update the flows. The latter approach is more stable and is thus what we employed in the final experiments. We provide more details in Appendix D. We note that the first training pipeline has the potential to allow for better exploration, and we therefore leave a more detailed exploration of it to future work.

### v3

@@ A.4 Parallelism versus Acceleration Time
Given two distributions $\pi^{n-1}$ and $\pi^{n}$, should we apply $K$-step forward and backward kernels; or insert K1K-1 intermediate distributions $(\pi^{n-1,k})_{k=1}^{K-1}$ and use only one-step forward and backward kernels instead?
By Proposition˜2, the inverse of the local round trip rate between $\pi^{n-1}$ and $\pi^{n}$ for the first method (time-accelerated) is
@@
The following proposition analyses these local rates for both methods. As in Section˜A.3, the main idea is to find a vanilla PT equivalent for both algorithms. We define
@@
This proposition shows that it is preferable to use the parallel-accelerated method, as the quantity in Equation˜17 is generally greater than that of Equation˜16 thanks to the effect of the bridge between ${\mathbb{S}}_{0}$ and ${\mathbb{S}}_{K}$. However, in practice the time-accelerated method consumes less memory and so might be more suitable in certain circumstances.
@@ Appendix B Further Details on Analysis of PT
B.1 Proof of Proposition˜2
Let $S^{n}_{t}$ be the indicator random variable with $S^{n}_{t}=1$ if the $n$-th swap is proposed and accepted at iteration $t$, and $S^{n}_{t}=0$ otherwise. By ˜1, the random variables $S^{n}_{t}$ are i.i.d. in $t$ with $S^{n}_{1},S^{n}_{2},\ldots\stackrel{{\scriptstyle d}}{{=}}\mathrm{Bernoulli}(s_{n})$, where $s_{n}:=1-r({\mathbb{P}}^{n-1},{\mathbb{Q}}^{n})$. This implies that the index processes $(I^{m}_{t},\epsilon^{m}_{t})$ form Markov chains on $\{0,\ldots,N\}\times\{-1,1\}$ with initial conditions $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ when $m$ is even, $\epsilon^{m}_{0}=1$ when $m$ is odd. For $t\geq 1$, each process satisfies the following recursion:
@@ B.1 Proof of Proposition˜2
$\epsilon^{m}_{t}$ $=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$ (19)
The remainder of the proof follows identically to (Syed et al., 2022, Corollary 2).
B.2 Proof of Proposition˜3
We first spell out the full statement of Proposition˜3.
@@ B.2 Proof of Proposition˜3
[…]
@@
where $r(p,q)=\|p\times q-q\times p\|_{\mathrm{TV}}$ and ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ are respectively the full path measures of Equation˜20 and Equation˜21.
@@
where the first inequality is elementary and the second follows from Proposition˜6. By data processing inequality
@@
Equation˜24 and Equation˜25 establish Equation˜23. On the other hand, putting
$R_{K}(x_{k})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{K}^{n}\times{\mathbb{P}}_{K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n})}(x_{k});$
$R_{\infty}(x_{\infty})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{\infty}^{n}\times{\mathbb{P}}_{\infty}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n})}(x_{\infty});$
$\varphi(\alpha)$ $:=1-\min(1,\alpha)$
we have by Lemma˜1
$r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})=\mathbb{E}_{{\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n}}[\varphi\circ R_{K}].$ (26)
Since $\varphi(\alpha)\in[0,1]$ the definition of the total variation distance implies
$\left|\mathbb{E}_{{\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n}}[\varphi\circ R_{K}]-\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]\right|\leq\mathcal{O}(1/\sqrt{K}).$ (27)
In addition let $D_{K}:\mathcal{C}[0,1]\to(\mathbb{R}^{d})^{K+1}$ be the discretisation operator which takes a continuous path $x_{\infty}$ and extract its values at $K+1$ discretisation points. Then
$\begin{split}\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]&=\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K}\circ D_{K}]\\ &\overset{K\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}]=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})\end{split}$ (28)
by dominated convergence theorem, taking into account the pathwise convergence of the Euler-Maruyama schemes (Kloeden & Neuenkirch (2007); Gyöngy (1998), see also Berner et al. (2025, Lemma B.7.a)). Equation˜26, Equation˜27, and Equation˜28 entail Equation˜22, finishing the proof.
@@
Consider the continuous-time extension $(\hat{X}_{t})_{t\in[0,1]}$ of the $K$-step Euler-Maruyama discretization of the SDE in Equation˜29 given by
@@
We use $\hat{{\mathbb{P}}}_{\infty}=\hat{{\mathbb{P}}}^{(K)}_{\infty}$ to denote its path-measure. By Pinsker’s inequality
@@
$\text{KL}({\mathbb{P}}^{*}_{K}\|{{\mathbb{P}}}_{K})\leq\text{KL}({{\mathbb{P}}}_{\infty}\|\hat{{\mathbb{P}}}_{\infty})=\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}\|b(t,X_{t})-b(t_{k},X_{t_{k}})\|^{2}dt.$
@@
$\mathbb{E}\|X_{t}-X_{t_{k}}\|^{2}$ $=\mathbb{E}\left\|\int_{t_{k}}^{t}b(s,X_{s})ds+\sigma(B_{t}-B_{t_{k}})\right\|^{2}$
$\leq 2(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}\left\|b(s,X_{s})\right\|^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 2G(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}(1+\|X_{s}\|)^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 4G(t-t_{k})\int_{t_{k}}^{t}(1+\mathbb{E}\|X_{s}\|^{2})ds+2d\sigma^{2}(t-t_{k})$
@@ B.3.1 Regularity Assumptions
Suppose $\pi^{\beta}(x)=\exp(-U^{\beta}(x))/Z_{\beta}$, where $Z_{\beta}=\int_{\mathcal{X}}\exp(-U^{\beta}(x))\mathrm{d}x$. For $\bm{\beta}=(\beta,\beta^{\prime})$, suppose ${\mathbb{P}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ and ${\mathbb{Q}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ are mutually absolutely continuous path measures on $\mathcal{X}^{K+1}$ with marginals ${\mathbb{P}}^{\bm{\beta}}(x_{0})=\pi^{\beta}(\mathrm{d}x_{0})$, and ${\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{K})=\pi^{\beta^{\prime}}(\mathrm{d}x_{K})$ with work functional $W^{\bm{\beta}}:\mathcal{X}^{K+1}\mapsto\mathbb{R}$,
@@
We will make the following regularity assumptions on ${\mathbb{P}}^{\bm{\beta}}$ and ${\mathbb{Q}}^{\bm{\beta}}$ analogous to Assumptions 5–7 in Syed et al. (2024). Suppose there exists $\mathcal{F}=(\mathcal{F}_{i})_{i\in\mathbb{N}}$ of a nested sequence $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots$ of vector spaces of measurable function $f:\mathcal{X}^{K+1}\to\mathbb{R}$, such that (1) $\mathcal{F}_{i}$ contains the constant functions, and (2) $\mathcal{F}_{i}$ is closed under domination, i.e., if $g$ is a measurable function such that $|g|\leq|f|$ for some $f\in\mathcal{F}_{i}$, then $g\in\mathcal{F}_{i}$.
@@
B.3.2 Proof of Theorem˜2
For $\bm{\beta}=(\beta,\beta^{\prime})$ recall the rejection rate can be expressed as $r({\mathbb{P}}^{\bm{\beta}},{\mathbb{Q}}^{\bm{\beta}})=1-{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\alpha^{\bm{\beta}}]$, where ${\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}$ is a measure over $\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}$ is the product between the forward and backwards paths,
@@ B.3.2 Proof of Theorem˜2
Suppose Assumptions 2–4 hold. For all $\bm{\beta}=(\beta,\beta^{\prime})$ with $\Delta\beta=\beta^{\prime}-\beta>0$, exists a constant $C>0$ independent of $\bm{\beta}$ such that,
@@
Given an annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$, let $\bm{\beta}_{n}=(\beta_{n-1},\beta_{n})$ and $\Delta\beta_{n}=\beta_{n}-\beta_{n-1}$. It follows from Lemma˜3 that the sum of the rejection satisfies,
@@
We will first approximate the first term in Equation˜30. Note that since $\Delta W^{\beta}=0$ and $\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}=\lim_{\beta^{\prime}\to\beta}\Delta W^{\bm{\beta}}/\Delta\beta=:\Delta\dot{W}^{\bm{\beta}}$, we have that $\dot{\alpha}^{\beta}$ satisfies:
@@
For the second term in Equation˜30, we have
@@
where $\bar{W}_{2}\oplus\bar{W}_{2}(x_{0:K},x_{0:K}^{\prime}):=\bar{W}_{2}(x_{0:K})+\bar{W}_{2}(x_{0:K}^{\prime})$. ˜2 guarantees that the expectation in the second-to-last line is continuous in $\bm{\beta}$, and hence $C_{2}$ is finite. Next, we note that since $|\Delta\dot{W}^{\beta}(x_{0:K},x_{0:K}^{\prime})|=|\Delta\dot{W}^{\beta}(x_{0:K}^{\prime},x_{0:K})|$ is symmetric and ${\mathbb{P}}^{\beta}={\mathbb{Q}}^{\beta}$, we have:
@@
By Lemma˜2, $\beta\mapsto\lambda_{\beta}$ is continuously differentiable. Since $\lambda_{\beta}\Delta\beta$ is a right Riemann sum for the integral of $\lambda_{\beta}$ with error:
@@ C.1.1 Work Formula for Deterministic flows
then, for all $(z^{n-1},z^{n})$ such that $T^{n}(z^{n-1})=z^{n}$, we have the following expression for the work defined in Equation˜3:
@@
and the acceptance rate $\alpha^{n}$ defined in Section˜3.2 becomes
@@
Similarly, applying Equation˜32 at $(z^{n-1},z^{n})=(x^{n-1}_{*},x^{n})\in S$ gives
@@
Together Equation˜33 and Equation˜34 establish the proposition. ∎
@@ C.1.2 Training
Since APT provides approximate samples from both the target and reference densities for each flow, it enables a range of training objectives. Some choices include maximum likelihood estimation (MLE) (equivalent to forward KL), reverse KL, and symmetric KL (SKL), which averages the two. Each has trade-offs: forward KL promotes mode-covering, reverse KL is more mode-seeking, and SKL balances both. APT’s parallel structure makes SKL particularly effective by providing access to samples at each intermediate annealing distribution, a feature many other methods lack. For example, sequential Monte Carlo (SMC)-based approaches such as FAB (Midgley et al., 2023) and CRAFT (Matthews et al., 2022) rely on samples from only one side, limiting their choice of loss functions.
Additionally, one can explore loss functions based on APT’s rejection rates, such as the analytic round trip rate, see Proposition˜2. Since higher round trip rates indicate more efficient mixing, optimizing for this metric improves sampling performance. Empirically, we found that using SKL yielded the most stable and robust results across different settings. This loss was therefore used in our final experiments. However, we leave it to future work to study other possible losses in more detail.
There are also possible variants to the training pipeline. In each case, we initialize normalizing flows to the identity transformation. One option is to run the APT algorithm with the current flows. After every (or several) steps of the APT algorithm, the current samples can be used to update the flow parameters. Since the flows are initialized at the identity transformation, the initial sampling of APT behaves similarly to PT. A second possible training pipeline is to instead use PT directly to generate a large batch of samples, and then use this batch of samples to update the flows. The latter approach is more stable and is thus what we employed in the final experiments. We provide more details in Appendix˜D. We note that the first training pipeline has the potential to allow for better exploration, and we therefore leave a more detailed exploration of it to future work.